On nostalgia night you decide to play Boardwalk. Remember that it is played with two six- sided dice. 1. You are in jail. If you roll doubles you get out. What are your odds of rolling doubles? You roll doubles. Now you are out. You get to roll again. What are your odds of rolling doubles on your second roll? Well it happened again. You rolled doubles a second time. What are your odds of rolling doubles on the third roll? What are your odds of rolling doubles three times in a row and going back to jail? Have you ever had to go to jail for three doubles? If you said yes, does it imply that you played monopoly a lot of times? a. b. c. d. e.
Answer #1

(a,b,c)

The odds of rolling a doubles on any roll, whether it be first roll or second roll or third roll is independent of previous rolls and thus these three odds have the same probability.

This probability is equal to 6/36 = 1/6

Since the sample space has 6*6=36 observations

And no of favourable outcomes are 6 which are

(1,1) (2,2) (3,3) (4,4) (5,5) (6,6)

Thus answers to a, b and c are 1/6 each.

(d)

The odds of getting doubles three times in a row is equal to

1/6 * 1/6 * 1/6

= 1/216

Thus answer to (d) is 1/216

(e)

Yes if I had to Jo to jail by getting doubles thrice in a row it implies that I play Monopoly. Because in every 256 rolls it is possible to get a doubles thrice in a row once